# Early warning signals of the tipping point in strongly interacting Rydberg atoms

### v1

@@ Results
Note here that the mean-field approximation is used and that the interaction terms ($V_{ij}^{r_{1}r_{2}}$, $V_{ij}^{r_{2}r_{1}}$, $V_{ij}^{r_{1}r_{1}}$, and $V_{ij}^{r_{2}r_{2}}$) are replaced by a single atom immersed in a field formed by these interactions between other atoms and this atom. Therefore, the task of solving the dynamic many-body problem is reduced to solving the dynamic problem for one single atom in that field. Besides, before the non-equilibrium phase transition occurs, the interaction between Rydberg atoms dissipates the atoms, which makes the dissipation ahead of the phase transition (the deviation from linear growth of the variance before the tipping point as shown in Fig. 1(e).). To model this process, we introduce a density-dependent dissipation in the mean-field approximation, and the detuning and the decay are replaced such that $\Delta_{1}\rightarrow\Delta_{1}-V\rho_{r_{1}r_{1}}^{\beta}$ and $\Gamma_{r_{1}}\rightarrow\Gamma_{r_{1}}-V\rho_{r_{1}r_{1}}^{\epsilon}$ [53, 23, 39, 10, 34], respectively, where $\beta$ and $\epsilon$ are exponent factors. By solving the master equation using the mean-field approximation and expanding the solution $\rho_{r_{1}r_{1}}$ as a function of $\Omega_{1}$ to the second order, the transmission against the intensity of the probe field $(\Omega_{1}/2\pi)^{2}$ is obtained. Here, $\beta=1$ and $\epsilon=3$ to enable better modeling of the results.
Next, we apply the MW field with dual tone frequencies to the theoretical model and simulate the transmission, where the MW field takes the form of $\Omega_{2}=E_{\mathrm{A}}(1+\mathrm{Cos}(\delta{f}t))$, and $E_{\mathrm{A}}$ is the magnitude of the Rabi frequency of the dual tone MW field. The simulated results for $\rho_{r_{1}r_{1}}$ versus $(\Omega_{1}/2\pi)^{2}$ are shown in Fig. 2(a-b), where panel (a) corresponds to the case of the phase transition without the MW field ($E_{\mathrm{A}}=0$) and panel (b) shows the case with the applied MW field ($E_{\mathrm{A}}\neq 0$). Figure 2(c) shows the variance that was calculated from Fig. 2(b), with reduced behavior in the region where 2.2 ($2.2\text{ }(\mathrm{MHz})^{2}<(\Omega_{1}/2\pi)^{2}<3.5\text{ }(\mathrm{MHz})^{2}$.
Figure 2: Theoretical and experimental results without and with the driving MW field. (a), (b) Theoretical simulations of the population $\rho_{r_{1}r_{1}}$ of the Rydberg atoms with different electric field amplitudes: (a) $E_{\mathrm{A}}$ = 0 $\mathrm{mV/cm}$ and (b) $E_{\mathrm{A}}\neq$ 0 $\mathrm{mV/cm}$. (c) Variance calculated from (b). (d), (e) Experimental transmission spectra under the different MW field amplitudes: (d) $E\rm_{MW}$ = 0 $\mathrm{mV/cm}$ and (e) $E\rm_{MW}$ = 1.64 $\mathrm{mV/cm}$. In this case, the scanning period Ts = 1 ms. (f) Measured variance from (e).
Transmission spectra under the dual-tone MW field driving conditions
By scanning the probe field intensity, we can observe the non-equilibrium phase transition in the Rydberg atom system, as illustrated in Fig. 2(d). To subtract the early warning signals before the tipping point is reached, we experimentally apply a dual-tone MW electric field to drive the transition between the two different Rydberg states $|50D_{5/2}\rangle$ and $|51P_{3/2}\rangle$, with a center frequency of $2\pi\times 17.041$ GHz and a frequency difference of $\delta f=2\pi\times 0.2$ MHz. The MW field induces a small ac Stark shift that moves the tipping point and varies the Rydberg atom population. The MW field causes the tipping point to move forward when compared with the case without the MW field, and the oscillation appears in the transmission spectrum in Fig. 2(e). By calculating the variance from the oscillated transmission, we observe a reduced variance when $(\Omega_{p}/2\pi)^{2}>26\text{ }(\mathrm{MHz})^{2}$, indicating the existence of an early warning signal, as shown in Fig. 2(f). The experimental results show good agreement with the expectations of the theoretical simulations.
From the results shown in Fig. 2(f), the measured variance is reduced nonlinearly when $(\Omega_{p}/2\pi)^{2}>27.5\text{ }(\mathrm{MHz})^{2}$. The increased $\Omega_{p}$ adds a number of excited Rydberg atoms, which makes the system dimension higher and then induces a strong critical slowing down effect. This occurs because, in a relatively low-dimensional system where the number of degrees of freedom is limited, localized interactions between neighboring Rydberg atoms are dominant. As a result, the critical slowing down process in a low-dimensional system is primarily influenced by these localized interactions. In this case, all the excited Rydberg atoms in the system can be synchronized rapidly, thus allowing faster recovery from the perturbation. In contrast, a high-dimensional system has large numbers of interacting Rydberg atoms, which results in a rich interconnectedness within the system. In this system, the critical slowing down effects are amplified because of the extended reach and the influence of the long-range interactions. Perturbations take longer times to propagate through the system, and the process of recovery from disturbances thus becomes slower.
Transmission spectra versus scanning time
The non-equilibrium phase transition in the Rydberg atom system can be modeled using a double-well transition, with each well standing for one bi-stable state, as depicted in Fig. 3(a)-(c). In the left well, phase 1 represents the few-body state, while in the right well, phase 2 corresponds to the many-body state. The dissipation of phase 2 is greater than that of phase 1 because of the interactions among the Rydberg atoms. When the population of Rydberg atoms is increased by $\Omega_{p}^{2}$, a phase transition occurs from the left well to the right well that crosses the barrier of the double well, as illustrated in Fig. 3(b). Near the tipping point $\Omega_{p,c}^{2}$, the critical slowing down phenomenon occurs; this term refers to an effect where the system’s response time to the driving MW field becomes slower (here, the MW field is regarded as a perturbation to the system). After the phase transition, the system is then in the I-phase, and the strong interactions between the Rydberg atoms make the system’s dissipation so large that the variance drops to a very low level.
A fast scan accumulates a small number of Rydberg atoms because of the small interval time. This causes the system to have low dissipation until the phase transition occurs, which means that the warning regime is almost invisible [see Fig. 3(d) and Fig. 3(e) for the scanning times of Ts= 0.25 ms and Ts= 0.5 ms, respectively]. As the light intensity increases, the system’s response to the external perturbation follows the increasing light intensity well.
@@
Figure 4: Trend of variance at different MW field intensities. (a) Variance measured at the different MW field intensities of $E_{\mathrm{MW}}$=1.80 mV/cm (blue), $E_{\mathrm{MW}}$=1.15 mV/cm (light green), and $E_{\mathrm{MW}}$=0.66 mV/cm (brown); see the Methods section for further details. To provide a better comparison of the signals at the different microwave intensities, we normalized the variance of the transmission for easier reading. The variance at 0.66 mV/cm is amplified by a factor of four and the variance at 1.15 mV/cm is amplified by a factor of 1.5, but the variance at 1.80 mV/cm is not amplified. (b) Theoretical simulations at the different modulation strengths of $E_{\mathrm{A}}$ = 2 (blue), $E_{\mathrm{A}}$ = 0.4 (black), and $E_{\mathrm{A}}$ = 0.1 (brown), where we normalized the simulated results using an amplification factor of 10 at $E_{\mathrm{A}}$ = 0.4 and an amplification factor of 20 at $E_{\mathrm{A}}$ = 0.1; the variance at $E_{\mathrm{A}}$ = 2 was not amplified. The red points correspond to the points where the variance diverged from the linear function in each case. In these processes, the tipping points for the different MW field intensities are calibrated to be in the same position for ease of comparison by slightly adjusting the coupling detuning.
@@
Figure 5: Variance measured at different temperatures. We measured the different variance characteristics at temperatures of (a) T=47.2C, (b) T=47.7C, (c) T=48.2C, and (d) T=49.1C. The results show that as the temperature increases, the divergence of the variance from the linear regime becomes more obvious. The early warning signal regimes shown in (a) and (d) are marked as the blue shaded areas for comparison. The solid red lines are the fittings with the function $y=p(x-20)e^{q(x-20)^{\alpha}}$+$s$, where $\alpha$ = 16.46 $\pm$ 0.22 for T=47.2C, $\alpha$ = 12.69$\pm$ 0.10 for T=47.7C, $\alpha$ = 10.89 $\pm$ 0.07 for T=48.2C, and $\alpha$ = 8.59 $\pm$ 0.07 for T=49.2C.
Transmission spectra at different temperatures
In our experiments, we also studied how the interaction between the Rydberg atoms affects the warning signals. A higher temperature leads to a higher interaction level, and vice versa. We changed the atomic density by varying the temperature and measured the variance of the oscillated transmission. Figure 5(a)-(d) show the measured results. We found that the variance before the phase transition was decreasing gradually as the temperature increased. For comparison, we found that the early warning signal regime shown in Fig. 5(d) was broader than that shown in Fig. 5(a), as given by ($\delta\Omega_{\mathrm{p}}/2\pi)^{2}=3.0\text{ }(\mathrm{MHz})^{2}$ at T = 47.2C and ($\delta\Omega_{\mathrm{p}}/2\pi)^{2}=1.5\text{ }(\mathrm{MHz})^{2}$ at T = 49.1C. Therefore, it is easier to identify the early warning signal at higher temperatures. This is explained by the fact that the interaction-induced dissipation causes the system to be difficult to synchronize when driven by the MW field, and it then leaves the linear response regime. This enhanced critical slowing down process in high-dimensional systems can be attributed to the increased system complexity and the interdependences among the Rydberg atoms.
The early warning signals of the Rydberg atom systems under different spatial structure conditions and scales [from their distinct densities] can also be probed, and the measured exponent $\alpha$ varies with the temperature [e.g., for T = 47.2C, $\alpha$ = 16.46 $\pm$ 0.22 in Fig. 5(a), but $\alpha$ = 8.59 $\pm$ 0.07 for T = 49.2C in Fig. 5(d)], thus indicating distinct critical dynamics before occurrence of a phase transition. This confirms that the early warning signal arises from the interactions between the Rydberg atoms, even when the system is still in the NI phase. This finding is consistent with the fact that the non-equilibrium phase transition in the Rydberg atom system occurs because of the interactions among a few local atoms, which eventually spreads to involve the atoms on a global scale. Using the MW field, we can detect small numbers of locally interacting Rydberg atoms, even without causing a phase transition to occur in the system.
Discussion
It should be noted that the measured variance in our experiment was reduced in the early warning regime, which differs from the increased variance that was reported previously [30, 31, 16]. Here, the variance is calculated from the transmission that occurs under the driving action of the external field, and it is proportional to the amplitude of the oscillation. In this way, the interaction-induced dissipation signal in the Rydberg atom system can be amplified, thus allowing us to characterize the full dynamics of the critical slowing down process. The early warning signals for the phase transition cause us to conclude that the interaction-induced dissipation appears first, and the phase transition occurs after that. This is also consistent with the results from our mean-field method, in which the exponent factor $\epsilon$ was greater than $\beta$.
Acknowledgements
We acknowledge funding from the National Key R&D Program of China (Grant No. 2022YFA1404002), the National Natural Science Foundation of China (Grant Nos. U20A20218, 61525504, and 61435011), the Anhui Initiative in Quantum Information Technologies (Grant No. AHY020200), the major science and technology projects in Anhui Province (Grant No. 202203a13010001)
Author contributions statement
D-S.D. conceived the idea. J.Z implemented the physical experiments. L-H.Z processed the microwave signal. D-S.D. and Z-K.L. conducted the theoretical model. D-S.D., J.Z, and Z-K.L. wrote the manuscript. All authors contributed to the experiment preparation and data discussions regarding the results. D-S.D. support this project.
Competing interests
The authors declare no competing interests.
Data and materials availability
All study data are included in the article.

### v2

@@ Abstract
Note here that the mean-field approximation is used and that the interaction terms ($V_{ij}^{r_{1}r_{2}}$, $V_{ij}^{r_{2}r_{1}}$, $V_{ij}^{r_{1}r_{1}}$, and $V_{ij}^{r_{2}r_{2}}$) are replaced by a single atom immersed in a field formed by these interactions between other atoms and this atom. Therefore, the task of solving the dynamic many-body problem is reduced to solving the dynamic problem for one single atom in that field. Besides, before the non-equilibrium phase transition occurs, the interaction between Rydberg atoms dissipates the atoms, which makes the dissipation ahead of the phase transition (the deviation from linear growth of the variance before the tipping point as shown in Fig.1(e).). To model this process, we introduce a density-dependent dissipation in the mean-field approximation, and the detuning and the decay are replaced such that $\Delta_{1}\rightarrow\Delta_{1}-V\rho_{r_{1}r_{1}}^{\beta}$ and $\Gamma_{r_{1}}\rightarrow\Gamma_{r_{1}}-V\rho_{r_{1}r_{1}}^{\epsilon}$ [23, 10, 34, 39, 56], respectively, where $\beta$ and $\epsilon$ are exponent factors. By solving the master equation using the mean-field approximation and expanding the solution $\rho_{r_{1}r_{1}}$ as a function of $\Omega_{1}$ to the second order. Here, $\Omega_{1}\propto\Omega_{p}\Omega_{c}/\Delta$, where $\Omega_{p}$ and $\Omega_{c}$ are Rabi frequencies of probe beam and coupling beam, respectively, and the $\Delta$ represents the detuning of intermediate state. This allows us to obtain the transmission against the intensity of the probe field [proportional to $\Omega_{p}^{2}$], which exhibits the same trend as $(\Omega_{1}/2\pi)^{2}$. Here, $\beta=1$ and $\epsilon=3$ to enable better modeling of the results.
@@
Figure 5: Trend of variance at different MW field intensities. (a) Variance measured at the different MW field intensities of $E_{\mathrm{MW}}$=1.80 mV/cm (blue), $E_{\mathrm{MW}}$=1.15 mV/cm (light green), and $E_{\mathrm{MW}}$=0.66 mV/cm (brown). To provide a better comparison of the signals at the different MW intensities, we normalized the variance of the transmission for easier reading. The variance at 0.66 mV/cm is amplified by a factor of four and the variance at 1.15 mV/cm is amplified by a factor of 1.5, but the variance at 1.80 mV/cm is not amplified. (b) Theoretical simulations at the different modulation strengths of $E_{\mathrm{MW}}$ = 1.8 mV/cm (blue), $E_{\mathrm{MW}}$ = 0.36 mV/cm (light green), and $E_{\mathrm{MW}}$ = 0.09 mV/cm (brown), where we normalized the simulated results using an amplification factor of 10 at $E_{\mathrm{MW}}$ = 0.36 mV/cm and an amplification factor of 20 at $E_{\mathrm{MW}}$ = 0.09 mV/cm; the variance at $E_{\mathrm{MW}}$ = 1.8 mV/cm was not amplified. The red points correspond to the points where the variance diverged from the linear function in each case. In these processes, the tipping points for the different MW field intensities are calibrated to be in the same position for ease of comparison by slightly adjusting the coupling detuning.
@@
Experimental setup
We used a two-photon transition scheme to excite the alkali metal atoms (rubidium 85) from the ground state to the Rydberg state. The probe beam ($\omega_{p}\approx$ 950 µm) is red-detuned by $2\pi\times$ 60 MHz, driving the atoms from the ground state $|g\rangle=|5S_{1/2},F=3\rangle$ to the intermediate excited state $|e\rangle=|5P_{3/2},F=4\rangle$, and the coupling beam ($\omega_{c}\approx$ 350 µm, $\Omega_{\mathrm{c}}/2\pi$ = 4.72 MHz) then drives the transition from $|e\rangle$ to the Rydberg state $|r_{1}\rangle=|50D_{5/2}\rangle$. We use a MW electric field to drive the RF transition between two different Rydberg states $|50D_{5/2}\rangle$ and $|51P_{3/2}\rangle$, and the MW electric field used in our experiment was generated by a RF source and a horn antenna. A 780 nm laser beam was first divided into two beams using a beam splitter, and the intensity of one of these beams was detected using a power meter. This method allows us to monitor the intensity of the experimental beam in real time. The other beam was split by a beam displacer (BD) into a probe beam and an identical reference beam, and both beams propagate in parallel through a heated Rb cell at a temperature of 47C. The probe beam is then overlapped with a counterpropagating coupling beam, thus constituting the EIT process. In the experiments, we caused the AOM to work in a linear range. We loaded a triangular wave signal that was generated using a signal generator (RIGOL DG4102) onto the AOM. In this way, we generated a probe beam with linearly increasing intensity ($(\Omega_{p}/2\pi)^{2}$ from 19.83 MHz2 to 35.22 MHz2). We used the Pound-Drever-Hall method to lock the probe field frequency, and the frequency of coupling beam is also fixed. By slightly adjusting the detuning of the coupling beam, we observed a sudden jump in probe transmission, which corresponds to the non-equilibrium transition from the phase 1 to the phase 2 for the Rydberg atom systems. When the dual-tone MW field was switched on, an oscillation effect occurs in the transmission spectrum. The transmission signals are detected using a differential photodetector and the data are collected through an oscilloscope.
Generation and calibration of the MW fields
The MW electric field used in our experiment was generated using an RF source (Ceyear 1465F-V) and a horn antenna. The RF source operates within the frequency range from DC to $2\pi\times 40$ GHz, and the horn antenna was located near the Rb cell. To calibrate the electric field amplitude $E_{\mathrm{MW}}$, we used the RF source to load an MW electric field signal with a frequency of $2\pi\times 17.041$ GHz, which is resonant with the transition of two Rydberg states, $|50D_{5/2}\rangle$ and $|51P_{3/2}\rangle$. Experimentally, we scanned the coupling detuning to obtain the EIT spectrum, and Autler-Townes splitting appeared in the EIT spectrum when the RF source was turned on. The MW electric field amplitude $E_{\mathrm{MW}}$ was calibrated based on the frequency interval $\Delta w$ of the Autler-Townes splitting.
$E_{\mathrm{MW}}=\frac{h\Delta w}{\mu_{d}}\cdot\frac{\lambda_{p}}{\lambda_{c}}$ (5)
where $h$ is Planck’s constant, $\mu_{d}$ is the transition dipole moment of the two Rydberg states, and $\lambda_{p}$ and $\lambda_{c}$ are the wavelengths of the probe field and the coupling field, respectively. To obtain the strength of the dual-tone MW field, we measured the signal from the RF source when connected directly to a spectrum analyzer (Ceyear 4024f) for calibration.